

********************************************************************
UARS WEEKLY STATUS REPORT - 24 JULY 1998
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation scan
data this past week without apparent problems.
***********
PEM
***********
PEM was not active.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
HALOE continues its operations and has been on since July 19. On about July 29
HALOE will turn off during a "graze" period and will be powered on again around
August 13 until August 20. The reprocessing of Version 19 HALOE data has been
underway and we have completed up to near the end of 1993, so far.
A concurrent Processing of Version 18 is continuing at Langley until the end of
1998, when the plan is to continue with Version 19 only.
***********
MLS
***********
MLS is scheduled to begin south viewing operations on Jul 24.
***********
HRDI
***********
HRDI is currently operating normally. We are scheduled to operate thru
July 23rd, then again from August 13th thru the 18th.
The time has come for the HRDI team to say goodbye and good luck to
Dave Ortland. Dave is leaving the beautiful climate of Michigan for
Northwest Research Associates in Seattle, Washington. Dave has been
with the HRDI group for over 10 years and has been responsible for
developing many of the HRDI inversion algorithms. All have been
challenging, but none more so than the retrieval of the stratosphere
wind. The successful development of this and other HRDI data products
would not have been possible without Dave's hard work, insight and
intuition, and attention to detail. Dave has performed many scientific
studies using HRDI data. His most recent work has contributed
significantly to our understanding of the wave inputs to the dynamics
of the tropical stratosphere. In addition, he has taught a formal
class on inversion theory and informal classes on atmospheric dynamics,
and chaired Ph.D. committees. Dave may be gone from the HRDI group,
but he will hardly be forgotten. He will continue to work with the
HRDI group on several outstanding problems. We wish him all the best
in his new research environs.
***********
WINDII
***********
WINDII WEEKLY REPORT FOR THE WEEK ENDING July 24, 1998
WINDII is currently off, and standing by for activation on August 8.
Team members are very much interested in the observation of OH Meinel
emission by ISAMS, reported in the July 1, 1998 issue of GRL. A contact
has been made to initiate comparison with WINDII OH Meinel observations.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 401 CD-ROM requests as of 24-July-1998.
There were six orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
86 days HAL (4.80) PROD (L1) (1873R-699C-15F=1159L)
90 days HAL (4.80) PROD (L3) (1873R-558C=1315L)
86 days SOL (4.62) PROD (L2-L3) (2349R-850C=1499L)
1 day SOL (4.62) TEST (1R-1C=0L)
1 day SUS (6.3) PROD (1R-1C=0L)
200 days SOL (4.62) PROD (L3) (200R-200C=0L)
1 day SOL (4.62) PROD (1R-1C=0L)
CDHF pending software upgrades: IDL 5.1, OpenIngres, Fortran V6.5.
***********
FOT
***********
REPORTING PERIOD
This report covers 11 July 1998 (Orbit 37322, GMT Day 192) through 17 July
1998 (Orbit 37426, GMT Day 198).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 192/2304 37336 Off
MLS 193/1008 37343 On/Off test
The beta angle reached a minimum angle of 0.0 degrees on 15 July 1998 (DOY
196). The beta angle is now increasing toward a maximum angle of 73.9
degrees on 05 Aug 1998 (DOY 217).
A Reverse-to-Forward Yaw Around maneuver (number 70) was successfully
executed in Orbit 37390 at 196/14:13:36 UTC on 15 July 1998. The SSPP
Instruments (ACRIM, SOLSTICE, and SUSIM) were enabled. Due to the SA
position at 269 degrees, the effective night length of the spacecraft after
the maneuver was 48 minutes. Post maneuver analysis of data indicate that
during the Yaw Around, the end-of-night load bus voltage was 23.8 volts and
the average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 23.0 and 22.5 percent respectively.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta between
Battery 2 and Battery 3 of 1.2 degrees C. Additional battery performance
data for each battery on 11 July 1998 and 17 July 1998 are provided in
Appendix D.
The clock error ranged from -10.3 to 9.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 06
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HALOE 01
% HRDI 03
% MLS 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/197 show the observatory data loss to be
20 hours, 45 minutes, 33.784 seconds (an increase of 32.096 seconds since
last report period). This is a 0.0347 percent data loss which equals a
99.9653 percent data capture for the mission.
The increase of 32.096 seconds occurred over two days:
* 0.608 seconds on Day 192 due to hits in the data, and
* 31.488 seconds on Day 193 due to data omission in dump.
Applications Processor (AP) ground system software testing has been
suspended waiting release of an operational pack containing version 10.11.
OTHER SIGNIFICANT EVENTS
Planning continues for the the UARS orbit adjust, scheduled for 12 August
1998, has been canceled. The orbit adjust will be re-scheduled in the
future at a time more consistent with reaching the 580 km lower operating
limit.
The B/D thruster refill effort continues. Planning continues and the
engineering thruster tests goal is still 28 July 1998.
Propulsion Module (PM) tank pseudos were added to the daily plots to show
amount of hydrazine fuel remaining. Also ran these plots at roughly
monthly intervals since just before the last orbit adjust (20 October
1994), in order to determine how much fuel was used and/or lost to leakage
during this time. Initial results seemed to indicate that about 4 pounds
of hydrazine had been inexplicably lost since the last orbit adjust. The
loss of fuel seemed to occur gradually over the period from late 1994 until
late 1996. However, further analysis showed that the apparent loss of fuel
was illusory, due to an indicated loss of fuel pressure, which resulted
from a proportional drop in average load bus voltage over the same time
span (fuel pressure telemetry calibration is a function of load bus
voltage). Based on this information, there is no indication of any
additional hydrazine fuel usage since the last orbit adjust, and we
currently estimate about 359 pounds of fuel remain.
The magnetic field model has been calibrated and aligned in the UARS Test
and Training Simulator (UTTS). Plans for testing On Board Computer (OBC)
earth pointing control (using the onboard magnetic software) are being
developed.
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
A list of representative telemetry points to be used as the test set for
verifying the decommutation subroutines in the re-engineered ground system
were identified. One orbit of CDHF data was processed to retrieve a
baseline test set of data, to be used to verify the new trending system
software.
FOT continues to support the UARS ground system phase-1 re-host effort to
the Lockheed Martin SCS-21 system. Work continued on the test script for
UARS TLM re-host phase-1 demonstration. Several meetings/telecons with the
SCS-21 personnel were supported to clarify UARS requirements for UARS re-
host.
MEETINGS
The FOT attended the UARS CMS Year 2K follow-up meeting on 16 July 1998.
At the meeting, the FOT agreed to assist Carey Myers and the CMS software
maintenance team in determining the status of the software releases which
currently reside on the back-up CMS system (UCMSB). The first step in
performing any CMS Year 2K testing is to get UCMSB up to speed with the
necessary CMS software and related support data. In addition, it was
decided at the meeting that the FOT should test and release the two pending
CMS software releases (R7B1V1 and R7B1V2) before any Year 2K related
software releases are developed and released
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 25 August
1998.
The UARS orbit adjust maneuver scheduled for 12 August 1998 has been
canceled and will be rescheduled at a later date.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 July 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37336 192/2304 98-239 HALOE Power OFF
37343 193/1008 98-240 MLS Test for Scan Mechanism
37352 194/0048 98-241 HRDI Heater Delay to 125 amp-min
37364 194/2010 98-242 HRDI Heater Delay to 150 amp-min
37368 195/0243 98-243 HRDI Heater Delay to 175 amp-min
37406 197/1507 98-244 OBC Clk Adj fr 0.00 to 20.6 ms/dy
37422 198/1651 98-245 OBC Clk Adj fr 20.6 to 0.00 ms/dy
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 July 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 37322-426
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 37327,41-3,72,76-
77,87,94,97-405,
417-18,21,24
NBRECBI Y-HI NBTR B Current 37377,94,417
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 37377,94
SOSTMPRNTRN Y-LO Temperature Entrance 37390-1,94,
97-403,17
ACTEMPANA2 Y-LO Temp Sensor Electronics 37394
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 37390-1 (YAW)
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 July 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37406 197/15:12 (16 Jul 98) from 00.0 to 20.6 msec/day
37422 198/16:57 (17 Jul 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 July 1998
11 Jul 1998 - GMT Day 192 Beta = 14.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.2 0.000 0.0
2 +16.8/ -39.2 +6.76/+5.47 +30.8/-14.4 24.2 1.021 23.1
3 +33.6/ -33.6 +5.47/+4.51 +29.6/-15.2 24.2 1.022 22.5
17 Jul 1998 - GMT Day 198 Beta = 6.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.2 0.000 0.0
2 +33.6/ -22.4 +5.47/+4.51 +29.6/-13.6 24.8 1.020 22.0
3 +44.8/ -5.6 +4.51/+2.92 +28.8/-15.2 24.6 1.020 21.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 July 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-044 FOT See list July 1998 14170 Generic Late Acquisition
SEVNTFW29